# Logistic-Beta Processes for Dependent Random Probabilities with Beta Marginals

### v2

@@ Abstract
The beta distribution serves as a canonical tool for modelling probabilities in statistics and machine learning. However, there is limited work on flexible and computationally convenient stochastic process extensions for modelling dependent random probabilities. We propose a novel stochastic process called the logistic-beta process, whose logistic transformation yields a stochastic process with common beta marginals. Logistic-beta processes can model dependence on both discrete and continuous domains, such as space or time, and have a flexible dependence structure through correlation kernels. Moreover, its normal variance-mean mixture representation leads to effective posterior inference algorithms. We illustrate the benefits through nonparametric binary regression and conditional density estimation examples, both in simulation studies and in a pregnancy outcome application.
@@ 1 Introduction
The beta distribution is widely used for modelling random probabilities owing to its flexibility, natural interpretation of parameters, and conjugacy in Bayesian inference. Beta distributions play a pivotal role in Bayesian nonparametric modelling (Hjort et al.,, 2010; Müller et al.,, 2015; Ghosal and van der Vaart,, 2017). For instance, a broad class of random probability measures have stick-breaking representations in terms of beta random variables, including the popular Dirichlet processes and Pitman-Yor processes (Ferguson,, 1973; Sethuraman,, 1994; Perman et al.,, 1992; Pitman and Yor,, 1997; Ishwaran and James,, 2001). Other notable models that build upon the beta distribution include Pólya trees (Ferguson,, 1974) for nonparametric density estimation and beta processes (Hjort,, 1990), which were originally designed as priors for the cumulative hazard rate but have been used broadly in machine learning due to their link with latent feature models, such as the Indian buffet process (Griffiths and Ghahramani,, 2011).
@@
Stochastic processes with beta marginals play a pivotal role in extending such models to handle non-exchangeable and heterogeneous data sources, while maintaining marginal properties; examples include dependent Dirichlet processes (MacEachern,, 1999; Quintana et al.,, 2022), dependent Pólya trees (Trippa et al.,, 2011) and dependent Indian buffet processes (Perrone et al.,, 2017). This article is motivated by the need to define flexible and computationally convenient classes of beta processes to substantially improve over current approaches, such as copula-based formulations (MacEachern,, 2000; Arbel et al.,, 2016; De Iorio et al.,, 2023). Copulas are highly flexible but lead to challenging posterior computation, particularly in complex settings. Trippa et al., (2011) and Bassetti et al., (2014) instead extend bivariate beta distributions (Olkin and Liu,, 2003; Nadarajah and Kotz,, 2005), but this can only induce positive correlation. Other strategies rely on latent binomial random variables (Nieto-Barajas et al.,, 2012) and transformation of Gaussian autoregressive processes (DeYoreo and Kottas,, 2018), but focused narrowly on discrete-time dependence. Finally, constructions based on covariate-dependent ordering (Griffin and Steel,, 2006), local sharing of random components (Chung and Dunson,, 2011), and Wright-Fisher diffusion (Perrone et al.,, 2017) are complex and limited to specialized cases.
@@
We propose a novel logistic-beta process, whose logistic transformation is a stochastic process with common beta marginals. The logistic-beta process builds upon the normal variance-mean mixture representation of a univariate logistic-beta distribution (Barndorff-Nielsen et al.,, 1982), and extends its hierarchical formulation to a stochastic process through correlation kernels in a related manner to a Gaussian process. As such, the proposed process can be flexibly defined on either discrete or continuous domains, and the implied dependence ranges from a perfect correlation to possibly negative correlations. Adopting the logistic-beta process prior for logit-transformed latent random probabilities provides a conditionally conjugate posterior sampling scheme that can directly exploit Pólya-Gamma data augmentation strategies (Polson et al.,, 2013).
@@
We provide two modeling examples utilizing the logistic-beta process: nonparametric regression for dependent binary data and conditional density estimation via a dependent Dirichlet process mixture with covariate-dependent weights and atoms. There are few methods for covariate-dependent weights in the dependent Dirichlet process due to computational challenges (Quintana et al.,, 2022; Wade et al.,, 2023). We illustrate how the proposed formulation leads to efficient posterior inference algorithms. Although there are alternative Bayesian nonparametric models for conditional densities (Dunson and Park,, 2008; Ren et al.,, 2011; Rigon and Durante,, 2021), maintaining a marginal Dirichlet process structure has advantages in terms of interpretability, ease of modelling dependence and alleviation of the hyperparameter sensitivity problem highlighted by Wade et al., (2023).
@@ 2.1 Univariate logistic-beta and Pólya mixing distributions
One important property of the logistic-beta distribution is its representation as a normal variance-mean mixture, as first shown by Barndorff-Nielsen et al., (1982). In particular, the density in equation (1) can be equivalently written as
@@
where $N_{d}(\mu,\Sigma)$ and $N_{d}(\eta;\mu,\Sigma)$ denote the distribution and density of a $d$-dimensional multivariate normal with mean $\mu$ and covariance $\Sigma$, and $\pi_{\textsc{po}}(\lambda;a,b)$ denotes the density of a Pólya distribution with shape parameters $a$ and $b$. This mixture representation is occasionally employed in regularized logistic regression problems (Gramacy and Polson,, 2012; Polson and Scott,, 2013). Figure 1 shows density plots of the logistic-beta and corresponding Pólya mixing distributions.
@@
See Appendix B for all proofs and additional details on the Pólya distribution.
@@ 2.2 Multivariate logistic-beta distributions
Definition 1.
Let $R$ be an $n\times n$ positive semidefinite correlation matrix. We say $\eta$ follows an $n$-dimensional multivariate logistic-beta distribution with shape parameters $a,b>0$ and correlation parameter $R$, denoted as $\eta\sim{\textsc{lb}}(a,b,R)$, if
$\eta\mid\lambda\sim N_{n}\left\{0.5\lambda(a-b)1_{n},\lambda R\right\},\quad\lambda\sim\mathrm{Polya}(a,b),$ (5)
where $1_{n}=(1,\dots,1)^{\mathrm{\scriptscriptstyle T}}\in\mathds{R}^{n}$.
Elementwise logistic transformation of a multivariate logistic-beta distribution induces a multivariate distribution on the unit hypercube $[0,1]^{n}$ with $\mathrm{Beta}(a,b)$ marginals. To see this, since $R$ has a unit diagonal, each component $\eta_{i}$ is marginally logistic-beta distributed with shape parameters $a$ and $b$ for $i=1,\dots,n$. Thus, logistic transformation $\{\sigma(\eta_{1}),\dots,\sigma(\eta_{n})\}^{\mathrm{\scriptscriptstyle T}}$ follows a multivariate beta distribution with $R$ controlling dependence. See Fig. 2 for illustrations. Such a multivariate generalization is mentioned in Barndorff-Nielsen et al., (1982), and the density is studied in Grigelionis, (2008). However, there is no literature using this construction for multivariate beta modeling. Alternative multivariate constructions that are related to the logistic-beta are proposed by Bradley et al., (2019) for conjugate spatio-temporal modeling of multinomial data and by Kowal et al., (2019) for temporal dependence in shrinkage.
@@
The main advantage of this construction is that the correlation parameter $R$ explicitly controls dependence with great flexibility. In Proposition 2 and Corollary 1, we present the first two moments of the multivariate logistic-beta and the range of pairwise correlation, illustrating how $R$ controls the strength of linear dependence.
Proposition 2.
Let $\eta\sim{\textsc{lb}}(a,b,R)$ with $R_{ij}$ being the $(i,j)$th element of $R$. Then, we have $E(\eta_{i})=\psi(a)-\psi(b)$, $\operatorname*{var}(\eta_{i})=\psi^{\prime}(a)+\psi^{\prime}(b)$ for $i=1,\dots,n$, and
@@
Corollary 1 (Correlation range).
Let $\eta\sim{\textsc{lb}}(a,b,R)$ with some $2\times 2$ positive semidefinite correlation matrix $R$. If $a=b$, $\operatorname*{corr}(\eta_{1},\eta_{2})$ has a full range $[-1,1]$, and if $a\neq b$, the range of $\operatorname*{corr}(\eta_{1},\eta_{2})$ is $[1-4\{\psi(a)-\psi(b)\}/[(a-b)\{\psi^{\prime}(a)+\psi^{\prime}(b)\}],1]$.
From Proposition 2, the covariance between $\eta_{i}$ and $\eta_{j}$ is a linear function of $R_{ij}$. The Pólya mixing variable $\lambda$ is shared across dimensions so that $R_{ij}=0$ does not imply independence between $\eta_{i}$ and $\eta_{j}$. The covariance between the logistic transformed variables $\operatorname*{cov}\{\sigma(\eta_{i}),\sigma(\eta_{j})\}$ lacks an analytical form, but Monte Carlo estimates are easily obtained. While Corollary 1 shows that arbitrarily high positive correlation can be captured for any choice of $(a,b)$, we have a nontrivial lower bound when $a\neq b$. The restriction on the strength of negative correlation is due to the shared asymmetric marginals (Joe,, 2006).
@@ 2.3 Logistic-beta processes and correlation kernels
Let $\mathcal{R}:\mathscr{X}\times\mathscr{X}\to[-1,1]$ be a correlation kernel which is positive semidefinite with $\mathcal{R}(x,x)=1$ for any $x\in\mathscr{X}$. We say $\{\eta(x)\in\mathds{R}:x\in\mathscr{X}\}$ is a logistic-beta process with shape parameters $a,b>0$ and correlation kernel $\mathcal{R}$, denoted as $\eta(\cdot)\sim{\textsc{lbp}}(a,b,\mathcal{R})$, if every finite collection $\{\eta(x_{1}),\dots,\eta(x_{n})\}^{\mathrm{\scriptscriptstyle T}}$ follows an $n$-dimensional multivariate logistic-beta with shape parameters $a,b$ and correlation parameter $R$ with $(i,j)$th element $R_{ij}=\mathcal{R}(x_{i},x_{j})$.
@@
From the definition of multivariate logistic-beta (5) that shares a common Pólya mixing variable and properties of the multivariate normal, the logistic-beta process is closed under marginalization and thus is a valid stochastic process as for other mixtures of Gaussian processes (Yu et al.,, 2007). Specifically when $a=b$, it belongs to the family of elliptical processes (Shah et al.,, 2014). The logistic transformation $\sigma\{\eta(\cdot)\}$ has marginals distributed as $\mathrm{Beta}(a,b)$, with dependence induced by the correlation kernel $\mathcal{R}$. Below we provide examples of correlation kernel constructions.
@@
The logistic-beta process can capture groupwise, temporal, and spatial dependence in discrete or continuous domains. When $\mathscr{X}$ corresponds to discrete time indices, we can consider an autoregressive order one ar(1) correlation kernel $\mathcal{R}(x,x^{\prime})=\rho^{|x-x^{\prime}|}$ for $|\rho|<1$. When $\mathscr{X}$ is a continuous spatial domain, a prominent example is the Matérn kernel $\mathcal{R}_{M}(x,x^{\prime};\varrho,\nu)=2^{1-\nu}\Gamma(\nu)^{-1}(\|x-x^{\prime}\|/\varrho)^{\nu}K_{\nu}(\|x-x^{\prime}\|/\varrho)$ with parameters $\varrho,\nu>0$, where $K_{\nu}$ is the modified Bessel function of the second kind; see Fig. 3 for realizations after logistic transformation.
@@
We can consider a more general construction of a correlation kernel $\mathcal{R}$. Let $\phi:\mathscr{X}\to\mathds{R}^{q}$ be a normalized feature map $\phi(x)=\{\phi_{1}(x),\dots,\phi_{q}(x)\}^{\mathrm{\scriptscriptstyle T}}$ such that $\|\phi(x)\|_{2}^{2}=1$ (Graf et al.,, 2003). We assume finite $q$, although infinite extensions are possible. Such a normalized feature map can be obtained from an unnormalized feature map $\varphi(x)$ with a bounded norm by letting $\phi(x)=\varphi(x)/\|\varphi(x)\|_{2}$; for example, $\varphi(x)$ can be spline basis functions for bounded domain $\mathscr{X}$, which induces nonlinear dependence in $x$. Then, a valid correlation kernel $\mathcal{R}$ can be defined as an inner product $\mathcal{R}(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle.$ Such kernels lead to convenient representations of the logistic-beta process presented in Proposition 3, which are helpful in understanding the induced dependence structure and facilitating posterior inference.
Proposition 3.
@@
Consider a normalized feature map $\phi:\mathscr{X}\to\mathds{R}^{q}$ and the corresponding kernel $\mathcal{R}(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle$. Let $\Phi=\{\phi_{k}(x_{i})\}_{ik}$ be an $n\times q$ basis matrix with $i$th row $\phi(x_{i})$. Then, a realization $\eta=\{\eta(x_{1}),\dots,\eta(x_{n})\}^{\mathrm{\scriptscriptstyle T}}$ from ${\textsc{lbp}}(a,b,\mathcal{R})$ can be equivalently written as the linear predictor representation (6) and the hierarchical representation (7),
$\eta$ =$=\{\psi(a)-\psi(b)\}(1_{n}-\Phi 1_{q})+\Phi\beta,\quad\beta\sim{\textsc{lb}}(a,b,I_{q}),$ (6)
$\eta$ $=0.5\lambda(a-b)1_{n}+\lambda^{1/2}\Phi\gamma,\quad\lambda\sim\mathrm{Polya}(a,b),\quad\gamma\sim N_{q}(0_{q},I_{q}).$ (7)
@@
Representation (6) resembles basis function representations of a Gaussian process with normal priors on coefficients (Gelman et al.,, 2015, §21.1), but with a $q$-dimensional multivariate logistic-beta prior on the coefficients. When we consider an independent Bernoulli response model with logit link $z(x)\sim\mathrm{Ber}[\sigma\{\eta(x)\}]$, equation (6) corresponds to the linear predictor of the logistic regression model with an additional fixed varying intercept term. The hierarchical model representation (7) is useful for posterior inference. Conditional on $\lambda$, the logistic-beta process with the normalized feature map kernel is parameterized with normal coefficients $\gamma$, enabling conditionally conjugate updates.

### v3

@@ Abstract
The beta distribution serves as a canonical tool for modeling probabilities in statistics and machine learning. However, there is limited work on flexible and computationally convenient stochastic process extensions for modeling dependent random probabilities. We propose a novel stochastic process called the logistic-beta process, whose logistic transformation yields a stochastic process with common beta marginals. Logistic-beta processes can model dependence on both discrete and continuous domains, such as space or time, and have a flexible dependence structure through correlation kernels. Moreover, its normal variance-mean mixture representation leads to effective posterior inference algorithms. We show how the proposed logistic-beta process can be used to design computationally tractable dependent Bayesian nonparametric models, including dependent Dirichlet processes and extensions. We illustrate the benefits through nonparametric binary regression and conditional density estimation examples, both in simulation studies and in a pregnancy outcome application.
@@ 1 Introduction
The beta distribution is widely used for modeling random probabilities owing to its flexibility, natural interpretation of parameters, and conjugacy in Bayesian inference. Beta distributions play a pivotal role in Bayesian nonparametric modeling (Hjort et al.,, 2010; Müller et al.,, 2015; Ghosal and van der Vaart,, 2017). For instance, a broad class of random probability measures have stick-breaking representations in terms of beta random variables, including the popular Dirichlet processes and Pitman-Yor processes (Ferguson,, 1973; Sethuraman,, 1994; Perman et al.,, 1992; Pitman and Yor,, 1997; Ishwaran and James,, 2001). Other notable models that build upon the beta distribution include Pólya trees (Ferguson,, 1974) for nonparametric density estimation and beta processes (Hjort,, 1990), which were originally designed as priors for the cumulative hazard rate but have been used broadly in machine learning due to their link with latent feature models, such as the Indian buffet process (Griffiths and Ghahramani,, 2011).
Stochastic processes with beta marginals play a pivotal role in extending such models to handle non-exchangeable and heterogeneous data sources. In the Bayesian nonparametric literature, a common solution lies in defining probability measure-valued stochastic processes whose realizations depend on covariates while maintaining marginal properties; examples include dependent Dirichlet processes (DDP; MacEachern,, 1999; Quintana et al.,, 2022), dependent Pólya trees (Trippa et al.,, 2011) and dependent Indian buffet processes (Perrone et al.,, 2017). Maintaining marginal properties is crucial since it not only simplifies how covariates influence the induced prior process but also clarifies the role of hyperparameters; see also MacEachern, (2000)’s four desiderata for dependent Bayesian nonparametric models.
@@
This article is motivated by the need to define flexible and computationally convenient classes of beta processes to substantially improve current approaches, such as copula-based formulations (MacEachern,, 2000; Arbel et al.,, 2016; De Iorio et al.,, 2023). Copulas are highly flexible but lead to challenging posterior computation, particularly in complex settings. Trippa et al., (2011) and Bassetti et al., (2014) instead extend bivariate beta distributions (Olkin and Liu,, 2003; Nadarajah and Kotz,, 2005), but this can only induce positive correlation. Other strategies rely on latent binomial random variables (Pitt et al.,, 2002; Pitt and Walker,, 2005; Nieto-Barajas et al.,, 2012) and the transformation of Gaussian autoregressive processes (DeYoreo and Kottas,, 2018), but focus narrowly on discrete-time dependence. Finally, constructions based on covariate-dependent ordering (Griffin and Steel,, 2006), local sharing of random components (Chung and Dunson,, 2011), and Wright-Fisher diffusion (Perrone et al.,, 2017) are complex and limited to specialized cases.
@@
We propose a novel logistic-beta process, whose logistic transformation is a stochastic process with common beta marginals. The logistic-beta process builds on the normal variance-mean mixture representation of a univariate logistic-beta distribution (Barndorff-Nielsen et al.,, 1982), and extends its hierarchical formulation to a stochastic process through correlation kernels in a manner related to a Gaussian process. As such, the proposed process can be flexibly defined in either discrete or continuous domains, and the implied dependence ranges from a perfect correlation to possibly negative correlations. Adopting the logistic-beta process prior for logit-transformed latent random probabilities provides a conditionally conjugate posterior sampling scheme that can directly exploit Pólya-Gamma data augmentation strategies (Polson et al.,, 2013).
@@
We provide two modeling examples utilizing the logistic-beta process: nonparametric regression for dependent binary data and conditional density estimation via a dependent Dirichlet process mixture with covariate-dependent weights and atoms. There are few methods for covariate-dependent weights in the dependent Dirichlet process due to computational challenges (Quintana et al.,, 2022; Wade and Inacio,, 2025). We illustrate how the proposed formulation leads to efficient posterior inference algorithms. Although there are alternative Bayesian nonparametric models for conditional densities (Dunson and Park,, 2008; Ren et al.,, 2011; Rigon and Durante,, 2021), maintaining a marginal Dirichlet process structure has advantages in terms of understanding the roles of parameters involved in the model, ensuring that marginal prior properties are not affected by covariates in the absence of such information, and alleviation of the hyperparameter sensitivity problem highlighted by Wade and Inacio, (2025).
@@ 2.1 Univariate logistic-beta and Pólya mixing distributions
An important property of the logistic-beta distribution is its representation as a normal variance-mean mixture, as first shown by Barndorff-Nielsen et al., (1982). In particular, the density in equation (1) can be equivalently written as
@@
where $N_{d}(\bm{\mu},\bm{\Sigma})$ and $N_{d}(\bm{\eta};\bm{\mu},\bm{\Sigma})$ denote the distribution and density of a $d$-dimensional multivariate normal with mean $\bm{\mu}$ and covariance $\bm{\Sigma}$, and $\pi_{\textsc{po}}(\lambda;a,b)$ denotes the density of a Pólya distribution with shape parameters $a$ and $b$. This mixture representation is occasionally employed in regularized logistic regression problems (Gramacy and Polson,, 2012; Polson and Scott,, 2013). Figure 1 shows density plots of the logistic-beta and corresponding Pólya mixing distributions.
@@
See Section S.1 of the Appendix for all proofs, and additional details on the Pólya distribution.
@@ 2.2 Multivariate logistic-beta distributions
Definition 2.1.
Let $\mathbf{R}$ be an $n\times n$ positive semidefinite correlation matrix. We say $\bm{\eta}$ follows an $n$-dimensional multivariate logistic-beta distribution with shape parameters $a,b>0$ and correlation parameter $\mathbf{R}$, denoted as $\bm{\eta}\sim{\textsc{lb}}(a,b,\mathbf{R})$, if
$\bm{\eta}\mid\lambda\sim N_{n}\left\{0.5\lambda(a-b)\bm{1}_{n},\lambda\mathbf{R}\right\},\quad\lambda\sim\mathrm{Polya}(a,b),$ (5)
where $\bm{1}_{n}=(1,\dots,1)^{\mathrm{\scriptscriptstyle T}}$.
Elementwise logistic transformation of a multivariate logistic-beta distribution induces a multivariate distribution on the unit hypercube $(0,1)^{n}$ with $\mathrm{Beta}(a,b)$ marginals. To see this, since $\mathbf{R}$ has a unit diagonal, each component $\eta_{i}$ is marginally logistic-beta distributed with shape parameters $a$ and $b$ for $i=1,\dots,n$. Thus, logistic transformation $\{\sigma(\eta_{1}),\dots,\sigma(\eta_{n})\}^{\mathrm{\scriptscriptstyle T}}$ follows a multivariate beta distribution with $\mathbf{R}$ controlling dependence. See Figure 2 for illustrations. Note that $\lambda\mathbf{R}$ in equation (5) is a valid covariance matrix since $\lambda>0$ and $\mathbf{R}$ is a positive semi-definite correlation matrix. Hence, the eigenvalues of $\lambda\mathbf{R}$ are simply the eigenvalues $\mathbf{R}$ scaled by $\lambda$. Such a multivariate generalization is mentioned in Barndorff-Nielsen et al., (1982), and the density is studied in Grigelionis, (2008). However, there is no literature using this construction for multivariate beta modeling. Alternative multivariate constructions related to logistic-beta are proposed by Bradley et al., (2019) for conjugate spatio-temporal modeling of multinomial data and by Kowal et al., (2019) for temporal dependence in shrinkage.
@@
The main advantage of this construction is that the correlation parameter $\mathbf{R}$ explicitly controls dependence with great flexibility. In Proposition 2.2 and Corollary 2.1, we present the first two moments of the multivariate logistic-beta and the range of pairwise correlation, illustrating how $\mathbf{R}$ controls the strength of linear dependence.
Proposition 2.2.
Let $\bm{\eta}\sim{\textsc{lb}}(a,b,\mathbf{R})$ with $R_{ij}$ being the $(i,j)$th element of $\mathbf{R}$. Then, we have $E(\eta_{i})=\psi(a)-\psi(b)$, $\operatorname*{var}(\eta_{i})=\psi^{\prime}(a)+\psi^{\prime}(b)$ for $i=1,\dots,n$, and
@@
Corollary 2.1 (Correlation range).
Let $\bm{\eta}\sim{\textsc{lb}}(a,b,\mathbf{R})$ with some $2\times 2$ positive semidefinite correlation matrix $\mathbf{R}$. If $a=b$, $\operatorname*{corr}(\eta_{1},\eta_{2})$ has a full range $[-1,1]$, and if $a\neq b$, the range of $\operatorname*{corr}(\eta_{1},\eta_{2})$ is $[1-4\{\psi(a)-\psi(b)\}/[(a-b)\{\psi^{\prime}(a)+\psi^{\prime}(b)\}],1]$.
From Proposition 2.2, the covariance between $\eta_{i}$ and $\eta_{j}$ is a linear function of $R_{ij}$, aiding interpretation of $\mathbf{R}$. The Pólya mixing variable $\lambda$ is shared across dimensions so that $R_{ij}=0$ does not imply independence between $\eta_{i}$ and $\eta_{j}$. The covariance between the logistic transformed variables $\operatorname*{cov}\{\sigma(\eta_{i}),\sigma(\eta_{j})\}$ lacks an analytical form, but Monte Carlo estimates are easily obtained. While Corollary 2.1 shows that arbitrarily high positive correlation can be captured for any choice of $(a,b)$, we have a nontrivial lower bound when $a\neq b$. The restriction on the strength of negative correlation is due to the shared asymmetric marginals (Joe,, 2006).
@@ 2.3 Logistic-beta processes and correlation kernels
Let $\mathcal{R}:\mathscr{X}\times\mathscr{X}\to[-1,1]$ be a correlation kernel which is positive semidefinite with $\mathcal{R}(x,x)=1$ for any $x\in\mathscr{X}$. We say $\{\eta(x)\in\mathds{R}:x\in\mathscr{X}\}$ is a logistic-beta process with shape parameters $a,b>0$ and correlation kernel $\mathcal{R}$, denoted as $\eta(x)\sim{\textsc{lbp}}(a,b,\mathcal{R})$, if every finite collection $\{\eta(x_{1}),\dots,\eta(x_{n})\}^{\mathrm{\scriptscriptstyle T}}$ follows an $n$-dimensional multivariate logistic-beta with shape parameters $a,b$ and correlation parameter $\mathbf{R}$ with $(i,j)$th element $R_{ij}=\mathcal{R}(x_{i},x_{j})$.
@@
From the definition of multivariate logistic-beta (5) that shares a common Pólya mixing variable and properties of the multivariate normal, the logistic-beta process is closed under marginalization and thus is a valid stochastic process as for other mixtures of Gaussian processes (Yu et al.,, 2007). Specifically when $a=b$, it belongs to the family of elliptical processes (Shah et al.,, 2014). The logistic transformation $\sigma\{\eta(x)\}$ has marginals distributed as $\mathrm{Beta}(a,b)$, with dependence induced by the correlation kernel $\mathcal{R}$. In the following, we provide examples of correlation kernel constructions.
@@
The logistic-beta process can capture groupwise, temporal, and spatial dependence in discrete or continuous domains. When $\mathscr{X}$ corresponds to discrete time indices, we can consider an autoregressive order one correlation kernel $\mathcal{R}(x,x^{\prime})=\rho^{|x-x^{\prime}|}$ for $|\rho|<1$. When $\mathscr{X}$ is a continuous spatial domain, a prominent example is the Matérn kernel $\mathcal{R}_{M}(x,x^{\prime};\varrho,\nu)=2^{1-\nu}\Gamma(\nu)^{-1}(\|x-x^{\prime}\|/\varrho)^{\nu}K_{\nu}(\|x-x^{\prime}\|/\varrho)$ with parameters $\varrho,\nu>0$, where $K_{\nu}$ is the modified Bessel function of the second kind; see Figure 3 for realizations after logistic transformation. In Figure S.1.2 of the Appendix, we illustrate how $\operatorname*{corr}\{\eta(x),\eta(x^{\prime})\}$ and $\operatorname*{corr}[\sigma\{\eta(x)\},\sigma\{\eta(x^{\prime})\}]$ change as a function of $\|x-x^{\prime}\|$ under Matérn correlation kernels and different choices of $(a,b)$.
Figure 3: Illustration of the logistic-beta process on $\mathscr{X}=[0,3]$ with Matérn correlation kernel. (Left) Eight realizations of $\sigma\{\eta(x)\}$ with $\eta(x)\sim{\textsc{lbp}}(2,4,\mathcal{R}_{M})$, following $\mathrm{Beta}(2,4)$ marginally. (Right) Posterior mean (solid) and 95% credible interval of latent success probabilities given the binary data $z(x_{i})\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{Ber}[\sigma\{\eta(x_{i})\}]$, $i=1,\dots,n$ represented as small vertical bars. The dashed line corresponds to the true success probabilities.
@@
We can consider a more general construction of a correlation kernel $\mathcal{R}$. Let $\bm{\phi}:\mathscr{X}\to\mathds{R}^{q}$ be a normalized feature map $\bm{\phi}(x)=\{\phi_{1}(x),\dots,\phi_{q}(x)\}^{\mathrm{\scriptscriptstyle T}}$ such that $\|\bm{\phi}(x)\|_{2}^{2}=1$ (Graf et al.,, 2003). We assume a finite $q$, although infinite extensions are possible. This normalized feature map can be obtained from an unnormalized feature map $\bm{\varphi}(x)$ with a bounded norm by letting $\bm{\phi}(x)=\bm{\varphi}(x)/\|\bm{\varphi}(x)\|_{2}$; for example, $\bm{\varphi}(x)$ can be spline basis functions for bounded domain $\mathscr{X}$, which induces non-linear dependence in $x$. Then, a valid correlation kernel $\mathcal{R}$ can be defined as an inner product $\mathcal{R}(x,x^{\prime})=\langle\bm{\phi}(x),\bm{\phi}(x^{\prime})\rangle.$ Such kernels lead to convenient representations of the logistic-beta process presented in Proposition 2.3, which are helpful in understanding the induced dependence structure and facilitating posterior inference.
Proposition 2.3.
@@
Consider a normalized feature map $\bm{\phi}:\mathscr{X}\to\mathds{R}^{q}$ and the corresponding kernel $\mathcal{R}(x,x^{\prime})=\langle\bm{\phi}(x),\bm{\phi}(x^{\prime})\rangle$. Let $\bm{\Phi}=\{\phi_{k}(x_{i})\}_{ik}$ be an $n\times q$ basis matrix with $i$th row $\bm{\phi}(x_{i})$. Then, a realization $\bm{\eta}=\{\eta(x_{1}),\dots,\eta(x_{n})\}^{\mathrm{\scriptscriptstyle T}}$ from ${\textsc{lbp}}(a,b,\mathcal{R})$ can be equivalently written as the linear predictor representation (6) and the hierarchical representation (7),
$\bm{\eta}$ =$=\{\psi(a)-\psi(b)\}(\bm{1}_{n}-\bm{\Phi}\bm{1}_{q})+\bm{\Phi}\bm{\beta},\quad\bm{\beta}\sim{\textsc{lb}}(a,b,\mathbf{I}_{q}),$ (6)
$\bm{\eta}$ $=0.5\lambda(a-b)\bm{1}_{n}+\lambda^{1/2}\bm{\Phi}\bm{\gamma},\quad\lambda\sim\mathrm{Polya}(a,b),\quad\bm{\gamma}\sim N_{q}(\bm{0}_{q},\mathbf{I}_{q}).$ (7)
@@
Representation (6) resembles basis function representations of a Gaussian process with normal priors on coefficients (Gelman et al.,, 2015, §21.1), but with a $q$-dimensional multivariate logistic-beta prior on the coefficients. When we consider an independent Bernoulli response model with logit link $z(x_{i})\sim\mathrm{Ber}[\sigma\{\eta(x_{i})\}]$, equation (6) corresponds to the linear predictor of the logistic regression model with an additional fixed varying intercept term. The hierarchical model representation (7) is useful for posterior inference. Conditional on $\lambda$, the logistic-beta process with the normalized feature map kernel is parameterized with normal coefficients $\bm{\gamma}$, enabling conditionally conjugate updates.